### v1

# An Improved Ensemble-Based Machine Learning Model with Feature Optimization for Early Diabetes Prediction

Abstract
Diabetes is a serious worldwide health issue, and successful intervention depends on early detection. However, overlapping risk factors and data asymmetry make prediction difficult. To use extensive health survey data to create a machine learning framework for diabetes classification that is both accurate and comprehensible, to produce results that will aid in clinical decision-making. Using the BRFSS dataset, we assessed a number of supervised learning techniques. SMOTE and Tomek Links were used to correct class imbalance. To improve prediction performance, both individual models and ensemble techniques such as stacking were investigated. The 2015 BRFSS dataset, which includes roughly 253,680 records with 22 numerical features, is used in this study. Strong ROC-AUC performance of approximately 0.96 was attained by the individual models Random Forest, XGBoost, CatBoost, and LightGBM.The stacking ensemble with XGBoost and KNN yielded the best overall results with 94.82% accuracy, ROC-AUC of 0.989, and PR-AUC of 0.991, indicating a favourable balance between recall and precision. In our study, we proposed and developed a React Native-based application with a Python Flask backend to support early diabetes prediction, providing users with an accessible and efficient health monitoring tool.
1 Introduction
Diabetes is a group of interconnected disorders that are characterized by high blood sugar. Diabetes occurs when the pancreas is unable to produce sufficient insulin or is unable to produce insulin at all, or the body cells fail to respond appropriately to insulin [1]. There are 2 type of diabetes Type 1 diabetes and Type 2 diabetes. Type 1 diabetes develops when the cells in the pancreas that secrete insulin are attacked and destroyed by the immune system [2]. Type 2 diabetes, on the contrary, develops when the body fails to respond to insulin (insulin-saturated fat and salt, age, chronic stress, high blood pressure, and low-density lipoprotein (LDL, ”bad”) cholesterol levels). Type 1 diabetes is more likely to be preceded by more environmental causes and infections in children than type 2 diabetes.A report by the World Health Organization (WHO), published in 2024, quotes that. In the year 2022, 830 million people were diagnosed with diabetes, which was 200 million in 1990 [2]. Fifty percent of the people suffering from diabetes are not being treated, which can lead to serious complications even now, such as blindness, kidney failure, heart attack, stroke, and amputation [3].
Diabetes is an emerging global health burden that needs early and accurate detection to reduce severe complications and improve outcomes [4]. Though numerous research efforts have applied machine learning to predict diabetes, many of these studies are based on limited or outdated datasets that cannot capture the complexity and diversity of real-world populations [5]. Such datasets are imbalance, noisy data, and overlapping health indicators, which reduce model reliability and generalization.
In this regard, the paper develops an optimized machine learning framework for the prediction of diabetes, using the BRFSS 2015 dataset with more than 250,000 . It aims at overcoming data imbalance and redundancy by advanced preprocessing, enhancing the prediction performance by using ensemble learning methods such as stacking, which could improve the accuracy, interpretability, and clinical utility of diabetes risk prediction models. Diabetes prediction using machine learning has also been done with the PIMA Indian and Vanderbilt datasets. The works follow the general pipeline of data preprocessing, feature selection, and classification using models like Logistic Regression, Random Forest, SVM, and ensembling techniques such as XGBoost and LightGBM with very high accuracy. Most of the models in existence are based on small or region-specific datasets, which do not allow for generalization into larger populations. Besides, challenges like data imbalance, noise, and low interpretability decrease the reliability of predictions. To overcome these gaps, this work uses the BRFSS 2015 dataset, applying SMOTE + Tomek Links and ensemble stacking techniques for building a more accurate, balanced, and interpretable diabetes prediction framework. The main aim of this study is to develop an accurate and efficient diabetes prediction model using advanced machine learning and ensemble techniques. In the proposed system, major improvements will be made at the data preprocessing stage, in feature selection, and in class balancing to enhance predictive performance. The main contributions of this work are listed as below:
  1. 1.
    Applied an advanced preprocessing pipeline that removes duplicates, imputes missing values, and normalizes data in order to prepare clean and reliable.
  2. 2.
    Applied SMOTE + Tomek Links for effective handling of data imbalance and noise reduction.
  3. 3.
    Applying the Ensemble Feature Selection Technique to rank the most important diabetes predictors among all the variables: MI, RFE, and LASSO.
  4. 4.
    A stacking ensemble model is developed where KNN was combined with XGBoost and was the best performer. The accuracy score it returns is high, 94.8%, with a ROC-AUC score of 0.989.
  5. 5.
    Built an interactive web application with a conversational chatbot for personalized health assistance and a prediction system where users input clinical factors to receive instant diabetes risk assessments. The platform provides doctors with feature importance visualizations to identify key risk factors, bridging the gap between advanced machine learning and practical clinical decision-making.
2 Methodology
This section explains the overall process used in our study, including data collection, preprocessing, feature selection, model development, and evaluation. Each step was carefully designed to ensure accurate, reliable, and interpretable results for diabetes prediction using machine learning.
2.1 Data Source and Data Preprocessing
The study used the 2015 BRFSS dataset containing 253,680 records with 22 numerical clinical features related to diabetes risk factors (age, glucose levels, BMI, blood pressure, etc.).
Class imbalance was addressed using SMOTE for oversampling minority cases and Tomek Links for undersampling majority cases. Features were normalized to standardize scale, and missing values were handled appropriately. The dataset was split into training (70%) and test (30%) sets for model evaluation.
Figure 1: Overall Workflow of Our Proposed Model
2.2 Analysis of Exploratory Data (EDA)
We conducted EDA to understand data distributions, patterns, and quality. Univariate analysis using histograms revealed class imbalance and outliers in continuous and categorical variables. Correlation analysis and VIF identified multicollinearity, with blood pressure, cholesterol, and BMI showing strong associations with diabetes, while other variables proved redundant. These findings guided subsequent feature selection to retain only predictive features.
2.3 Feature Selection
Three feature selection techniques were used to increase efficiency and decrease dimensionality: LASSO regression, Recursive Feature Elimination (RFE), and Mutual Information (MI). A single table was created by normalising the ranks derived from different techniques. The top 18 predictors were chosen for model training based on the average ranking: [’GenHlth’, ’HighBP’, ’BMI’, ’Age’, ’HighChol’, ’CholCheck’, ’Income’, ’Sex’, ’HeartDiseaseorAttack’, ’HvyAlcoholConsump’, ’AnyHealthcare’, ’DiffWalk’, ’PhysActivity’, ’Smoker’, ’Veggies’, ’Fruits’, ’Education’, ’Stroke’]. This choice lowers the chance of overfitting while striking a balance between explainability and predictive power.
2.4 Data Splitting
To maintain a consistent class distribution, stratified sampling was employed to split the dataset into 80% training and 20% testing. The test data was set aside for the final assessment to ensure that the findings were objective, while the training data was used to construct models and adjust hyperparameters using k-fold cross-validation.
2.5 Model Training Strategy
We trained nine supervised learning algorithms ranging from simple models (Logistic Regression, Naive Bayes) to advanced ensembles (Random Forest, XGBoost, LightGBM, CatBoost) to compare explainability against predictive power.
Hyperparameter optimization used a two-step approach: Randomized Search to explore the parameter space broadly, followed by GridSearchCV to fine-tune promising regions for optimal performance. We applied 5-fold cross-validation during training to prevent overfitting and ensure robustness. Final evaluation on the test set used multiple metrics (accuracy, precision, recall, F1-score, ROC-AUC, confusion matrix) for comprehensive model assessment.
2.6 Development of the Stacking Ensemble Model
We chose to use stacking instead of relying on a single model. This approach combines multiple algorithms to leverage their individual strengths. Looking at Table 1, XGBoost caught our attention. It delivered balanced, strong performance across accuracy, precision, and F1-score. But it had a weakness: recall was relatively low. Why does recall matter? In medical applications, missing even one diabetes case is problematic. We need to identify all positive cases. KNN took a different approach. It excelled at recall—meaning it caught diabetic patients really well. The downside was slightly lower overall accuracy. We realized these two models complemented each other perfectly. So we stacked them together using LightGBM as the meta-learner.
This combination worked well. By blending XGBoost’s balanced performance with KNN’s strong recall, we built a more robust model. It caught more positive cases while maintaining solid accuracy.
3 Results Analysis
This section presents and discusses the outcomes of our experiments, including the setup, model evaluation, and comparison with previous studies. The goal is to analyze how each model performed, interpret key findings from the results, and highlight the effectiveness of the proposed stacking ensemble model for diabetes prediction.
3.1 Performance Analysis of All Models
We tested nine machine learning models to find the best approach for diabetes prediction from clinical features. This included traditional methods like Logistic Regression and Naive Bayes, as well as advanced ensemble techniques like Random Forest, XGBoost, LightGBM, and CatBoost.
The stacking ensemble delivered the strongest results, achieving 94.82% accuracy on test data by combining the power of two complementary models. KNN provided exceptional sensitivity (capturing 97.02% of diabetic cases), while XGBoost contributed high precision (93% of positive predictions were correct). By blending these strengths, the stacking approach achieved the best balance across all performance metrics—particularly the highest F1-score of 0.9493, which matters most in medical applications. Naive Bayes was the weakest model with only 71.29% accuracy. Its core weakness is assuming that clinical features are independent when they’re actually interconnected (glucose levels, body weight, and age all influence each other). This fundamental limitation prevented it from capturing the complex patterns needed for accurate diabetes prediction.
Table 1: Performance Metrics of Applied Models for Early Diabetes Prediction
Model Accuracy Precision Recall F1 ROC-AUC
LightGBM 0.9108 0.9694 0.8483 0.9048 0.9651
Random Forest 0.9096 0.9389 0.8762 0.9064 0.9678
CatBoost 0.9074 0.9596 0.8506 0.9018 0.9619
Logistic Regression 0.7386 0.7257 0.7672 0.7459 0.8135
Naive Bayes 0.7129 0.6683 0.8453 0.7464 0.7750
XGBoost 0.9063 0.9642 0.8440 0.9001 0.9628
KNN 0.9023 0.8635 0.9557 0.9072 0.9417
Stacking 0.9482 0.9293 0.9702 0.9493 0.9895
Table 1, shows how all nine models performed. The standout finding: ensemble methods significantly outperformed traditional models. Tree-based models (Random Forest, XGBoost, LightGBM, CatBoost) clustered around 90-91% accuracy, while simpler statistical methods lagged below 74%.
Table 2: Training and Testing Accuracy Comparison of Applied Models
Model Train Accuracy Test Accuracy
Stacking 0.9875 0.9482
LightGBM 0.9185 0.9100
Random Forest 0.9584 0.9095
CatBoost 0.9396 0.9072
XGBoost 0.9203 0.9056
KNN 0.9919 0.8996
Logistic Regression 0.7413 0.7399
Linear SVC 0.7398 0.7369
Naive Bayes 0.7144 0.7129
Table 2, reveals a critical insight: the stacking model maintains excellent generalization, with only a 3.93% drop from training to test accuracy. In contrast, KNN shows severe overfitting, losing 9.23% performance on unseen data—a key reason why individual KNN couldn’t be relied upon. Ensemble methods consistently maintained stable performance gaps, demonstrating superior generalization compared to traditional statistical approaches.
3.2 Performance Measure for Ensemble Model
Figure 2: Performance of the Stacking Model: Confusion Matrix
Figure 2, shows the stacking model achieved 36,041 true negatives and 37,589 true positives, with minimal misclassifications (2,821 false positives and 1,273 false negatives), demonstrating balanced diagnostic performance across both classes.
Figure 3: Comparison of ROC Curves for All Models
The ROC curve Figure 3, evaluates how well each model distinguishes between diabetic and non-diabetic patients. The stacking ensemble dominates with an AUC of 0.989, meaning it correctly separates the two classes nearly perfectly across all decision thresholds. Random Forest (AUC = 0.968) and LightGBM (AUC = 0.965) also perform strongly, while traditional methods like Naive Bayes (AUC = 0.775) and Logistic Regression (AUC = 0.814) lag significantly behind. This clear separation shows why ensemble methods are superior for clinical diagnostic tasks.
Figure 4: Comparison of Precision–Recall Curves for All Models
The Precision-Recall curve Figure 7, tells the real-world story for medical applications where catching diabetic cases (recall) while minimizing false alarms (precision) both matter equally. The stacking model achieves an average precision of 0.991, the highest among all models. This exceptional performance means the stacking ensemble reliably identifies diabetic patients while maintaining high confidence in its positive predictions. Random Forest (AP = 0.974) and LightGBM (AP = 0.973) follow closely, but the stacking approach’s 0.991 score confirms its superior reliability for clinical deployment.
3.3 Comparison with Existing Works
Our proposed method outperforms previous work in terms of dataset size, hyperparameter tuning, feature selection, and imbalance handling. While some papers proposed slightly higher accuracy in different datasets with small numbers of data, our proposed model is highly competitive with large datasets (253,680 patients). Two papers proposed somewhat worse performance than our model: they achieved 86.6% and 92.5% accuracy, while we achieved 94.82% accuracy with a better AUC-ROC (0.9895). Table 3 contains a comprehensive summary of prior work.
Table 3: Performance comparison of proposed stacking model with previous studies in terms of dataset size, dataset source, and accuracy.
Reference Dataset Method Accuracy (%)
Rahaman [6] 768 (PIMA) Random Forest 81
Gupta [7] 768 (PIMA) Logistic Regression 86
Ha [8] PIMA + hospital + wearables Random Forest 90
El Massari [9] 768 (PIMA) Ontology Classifier 77.5
Khokhar [10] 253,680 (BRFSS) Ensemble model 92.5
Ren [11] 253,680 (BRFSS) CatBoost Classifier 86.6
Our Model 253,680 (BRFSS) KNN + XGBoost 94.82
In comparison to previous studies, our model achieves a good balance of high prediction performance and scalability.
3.4 Deployment Interface
Figure 5: User-provided clinical factors and model prediction results.
Figure 6: SugarSense chatbot providing real-time health assistance.
Figure 7: Feature importance visualization for key risk factors.
In this study, a user-friendly web-based application was developed to demonstrate the practical implementation of the proposed diabetes prediction model. The application allows users to input personal and health-related data such as age, gender, BMI, blood pressure, glucose level, and insulin measurements. After submission, the backend machine learning model analyzes the input data in real time and predicts whether the individual is diabetic or non-diabetic. As illustrated in Figure 5,6, and 7, the web interface displays the prediction outcome along with a confidence score, ensuring result transparency and interpretability.
4 Discussion
In this study, our model has achieved the highest accuracy of 94.82%, which is better than previous works. Moreover, we achieved an ROC-AUC of 0.98 and a PR-AUC of 0.991. For handling the imbalanced data, we used SMOTE and Tomek Links. We implemented multiple supervised learning algorithms, including logistic regression, random forest, gradient boosting, XGBoost, LightGBM, CatBoost, KNN, and SVM, in developing predictive models for diabetes, and among the tested algorithms, the ensemble with XGBoost and KNN provided the best output. In conclusion, we developed a web-based application that will help healthcare professionals with diabetes prediction.
The BRFSS dataset is a survey-based dataset, which has reporting bias and missing entries. The data is highly imbalanced. Although we used SMOTE and Tomek Links to improve the dataset, the oversampling may not be perfectly reflective. This database, from 2015, is based on only one population. Performance may vary for different communities. The combination of datasets must be reliable. In the future, we can use cost-sensitive learning or adaptive boosting for imbalanced data instead of relying solely on SMOTE and Tomek Links and test this model on different populations or ethnic groups.
5 Conclusion
This study demonstrates the prediction of diabetes using the 2015 BRFSS dataset, which contains 253,680 records with 22 numerical features. SMOTE and Tomek Links were used to rectify the imbalanced data. We explored supervised learning algorithms, including logistic regression, random forest, gradient boosting, XGBoost, LightGBM, CatBoost, KNN, and SVM, in developing predictive models for diabetes, and among the tested algorithms, the ensemble with XGBoost and KNN provided the best accuracy of 94.82%, ROC-AUC of 0.989, and PR-AUC of 0.991. We developed an Web-based application.
This study shows that using hybrid feature selection with balanced data improves the diabetes classification performance. The data is very imbalanced. Although we used SMOTE and Tomek Links to improve the dataset, the oversampling may not be perfectly reflective. The BRFSS dataset is a survey-based dataset, which has reporting bias and missing entries. This database, from 2015, is based on a population. Performance may vary for different populations or ethnic groups. The combination of datasets must be reliable. For ensuring global applicability, the model may be tested on other countries and different ethnic groups, and may use cost-sensitive learning or adaptive boosting for imbalanced data instead of relying solely on SMOTE and Tomek Links.